when N is divided by 5, the remainder is 3. what is the remainder when 7N is divided by 5?
Answer by richwmiller(17219)

n/5=x+3
7n/5=y
n=5(x+3)
n=5x+15
7*(5x+15)/5
7*(x+3)
7x+21
21/7=3 no remainder
7x/7=x no remainder
So there is no remainder